### v3

@@ Abstract
The performance of quantum key distribution (QKD) heavily depends on statistical inference. For a broad class of protocols, the central statistical task is a random sampling problem, customarily addressed using exponential tail bounds on the hypergeometric distribution. Here we devise a strikingly simple exponential bound for this task, of unprecedented tightness among QKD security analyses. As a by-product, confidence intervals for the average of non-identical Bernoulli parameters follow too. These naturally fit in statistical analyses of decoy-state QKD and also outperform standard tools. Lastly, we show that, in a vast parameter regime, the use of tail bounds is not enforced because the cumulative mass function of the hypergeometric distribution is accurately computable. This sharply decreases the minimum block sizes necessary for QKD, and reveals the tightness of our simple analytical bounds when moderate-to-large blocks are considered.
Introduction. Quantum key distribution (QKD) is the only approach to key exchange that is provably secure against computationally unbounded adversaries [1]. Today, considerable efforts are being made to boost the performance of practical QKD systems while guaranteeing their physical-layer security [2, 3]. Among the multiple facets of this ongoing program, it is of utmost importance to achieve tighter lower bounds on the extractable secret key length, a goal that can be tackled by refining QKD security proofs. These proofs build upon quantum information-theoretic results and strongly rely on statistical inference, which plays a pivotal role to keep the necessary data block sizes to a minimum.
To be precise, establishing the finite-key security of a QKD protocol requires to upper bound the probability of a “failure” happening in its parameter estimation (PE) step. This refers to the event where the protocol does not abort but the adversary obtains non-trivial information about the final key [1]. In BB84-like protocols [4], a very popular proof technique is based on an entropic uncertainty relation (EUR) for smooth entropies [5, 6], in which case the simplest formulation of the problem goes as follows. Let us consider a two-piece experiment where a population of $N$ binary random variables is generated according to an arbitrary probability distribution, and a random sample of size $n$ is drawn from the population afterwards. If we denote the frequency of ones in the test (complementary) sample as $\hat{p}$ ($\hat{q}$), the goal is to find a threshold value $q^{\rm th}\geq{}p^{\rm th}$ such that
@@
for arbitrary values of $N$, $n$, $\epsilon$ and $p^{\rm th}$. Naturally, an interplay exists between $\epsilon$ and $q^{\rm th}$: fulfilling Eq. (1) with a very small $\epsilon$ comes at the price of tolerating larger values of $q^{\rm th}$. In the context of QKD —in which $\hat{q}$ is often referred to as the phase-error rate (PHER)— this translates into a direct penalty on the provably secure key length. Notably, the above formulation of the problem shows up exactly in ideal implementations of the BB84 [4] or the BBM92 [7, 8] protocols, and we adhere to the latter in the main text for didactic purposes. The detailed protocol description and the secret key length formula can be found in Appendix I. On the other hand, the popular decoy-state BB84 protocol [9, 10, 11, 12], which is implemented in most of today’s commercial QKD setups, is addressed in Appendix II.
@@
In EUR-based security proofs [8, 13], it is commonplace to derive the desired $q^{\rm th}$ by resorting to an exponential tail bound on the hypergeometric (HG) distribution known as Serfling inequality [14]. Recently though, a sharper bound has been derived in [15], combining Serfling inequality with a refined inequality due to Hush and Scovel [16]. Nonetheless, the attainable improvement is moderate and heavily relies on the use of very large samples for testing, an undesirable feature given the fact that the final key is extracted from the complementary data.
@@
In this paper, we propose a simpler approach to calculate $q^{\rm th}$, resorting to a well-known link between random sampling with and without replacement [17]. Remarkably, the bound we obtain outperforms [15] in all the explored parameter regimes relevant for QKD. Furthermore, in the derivation of the bound, confidence intervals (CIs) are obtained for the average parameter of a set of independent Bernoulli trials. These intervals, which provide an additional tool required for current decoy-state QKD security proofs, also outperform the analytical results considered for that purpose so far [12, 18]. Lastly, we show that, contrary to general belief, the cumulative mass function (CMF) of the HG distribution is exactly computable in a vast parameter regime of interest to QKD. Crucially, this enables the calculation of optimal —so-called Clopper-Pearson— one-sided CIs for the population parameter [19]. Applying the latter to the random sampling problem of Eq. (1) —and also to its generalization to decoy-state QKD schemes— we report a dramatic decrease of the minimum block sizes necessary for key extraction. On the other hand, explicit comparison with our simple analytical tools reveals the tightness of the latter for moderate-to-large blocks.
@@
Analytical results. Our analytical results are triggered by two observations. First, the additive Chernoff bound [20], conventionally used in the context of sampling with replacement, also holds for sampling without replacement. This feature, which is a consequence of a more general theorem by Hoeffding [17] —see also [21] for a simple proof— has already been exploited in QKD security analyses before [22, 23]. Secondly, regardless of the context, very loose relaxations of the bound are often used in the literature. With this in mind, we derive a tight relaxation of the additive Chernoff bound and indistinctly use it to obtain one-sided CIs for the population parameters in sampling with and without replacement. Then, we apply these CIs to address the problem of Eq. (1) and its generalization to decoy-state QKD.
@@
A proof of this claim is given in Appendix III. Importantly, since the claim is solely derived from the additive Chernoff bound, it automatically extends to a HG random variable [17, 21]. The upper extension goes as follows.
@@
As shown next, Proposition 2 allows to obtain a threshold value $q^{\rm th}$ fulfilling Eq. (1), in so providing the failure probability estimation of the BBM92 protocol.
@@
for the threshold function $q^{\rm th}_{N,n,\epsilon}(x)$ defined in Eq. (6), given that the latter is locally non-decreasing at $p^{\rm th}$.
@@
In the first inequality we use the fact that $\{\hat{p}\leq{}p^{\rm th},\hat{q}\geq{}q^{\rm th}_{N,n,\epsilon}\left(p^{\rm th}\right)\}\implies{}\{\hat{q}\geq{}q^{\rm th}_{N,n,\epsilon}\left(\hat{p}\right)\}$. This is so because $q^{\rm th}_{N,n,\epsilon}(x)$ is non-decreasing for $x\in[0,p^{\rm th}]$ if it is non-decreasing at $x=p^{\rm th}$, and the latter is guaranteed by assumption. The averaging over $\hat{T}$ in the second inequality is a purely ad hoc step aimed at enabling the use of Proposition 2, which presumes a fixed frequency of ones in the population. The third inequality follows from Proposition 2. ∎
A relevant observation is that the condition on the slope at $p^{\rm th}$, which is equivalent to $\Gamma^{+^{\prime}}_{n,\epsilon}(p^{\rm th})\geq{}n/N$ ($\Gamma^{+^{\prime}}_{n,\epsilon}(p^{\rm th})$ denoting the derivative of $\Gamma^{+}_{n,\epsilon}(x)$ with respect to $x$ evaluated in $p^{\rm th}$), is unrestrictive for all practical purposes [25].
@@
Numerical results. Certainly, tighter results can be obtained by resorting to numerical methods. For instance, using Stirling’s formula and the Gaussian cumulative density function, the authors of [22] construct a HG tail inequality that outpeforms our relaxed Chernoff bound in certain regimes, but does not enable a closed formula for the related CI. Alternatively, one could use the exact additive Chernoff bound rather than a relaxation. Note, however, that if one is willing to use numerical techniques, tail inequalities may become superfluous. This is so because, contrary to general belief, accurately computing the CMF of the HG distribution is often feasible in QKD applications. Crucially, this enables the calculation of the strictest possible one-sided CIs [26, 27] using the well-known Clopper-Pearson method based on pivoting the CMF [19]. In particular, for our purposes, one can replace $\Gamma^{+}_{n,\epsilon}(\hat{p})$ in Corollary 1 by the tighter statistic $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$, defined via [27]
@@
if the target set of the minimization is non-empty, and $\mathcal{CP}^{+}_{N,n,\epsilon}(x)=(N+1)/N$ otherwise [28], where we recall that $p$ takes discrete values in the set $\{0,1/N,2/N,\ldots{}1\}$ [29]. Remarkably, accurate determination of the CMF is required in order to compute $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$, with a numerical precision beyond $\epsilon$ according to Eq. (9). Indeed, for a broad range of settings of interest to QKD, this is a manageable task with existing software.
@@
Importantly, since $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$ fulfills Corollary 1, it also fulfills Proposition 2, and monotonicity considerations allow to further establish Proposition 3 replacing Eq. (6) by
$q^{\rm th}_{N,n,\epsilon}(x)=\frac{N\mathcal{CP}^{+}_{N,n,\epsilon}(x)-nx}{N-n}.$ (10)
@@
In Fig. 1 we compare the candidate threshold values $q^{\rm th}$ satisfying Eq. (1): Eq. (6), Eq. (10), and the thresholds derived in [15] and [8]. For illustration purposes, we consider typical QKD inputs, $p^{\rm th}\leq{}4\%$, $N=10^{5}$, $n=10^{4}$ and $\epsilon=10^{-9}$, and remark that the improvement of our tools —i.e. Eq. (6) and Eq. (10)— prevails in all the explored parameter regimes relevant for QKD.
@@
Figure 1: Threshold value $q^{\rm th}$ satisfying Eq. (1), for $p^{\rm th}\leq{}4\%$, $N=10^{5}$, $n=10^{4}$ and $\epsilon=10^{-9}$. The solid red (blue) line corresponds to our numerical (analytical) tool, based on the exact cumulative mass function of the hypergeometric distribution (a relaxed additive Chernoff bound [20]). The dashed orange line uses the result in [15], which combines Serfling [14] and Hush & Scovel [16] inequalities, and the dotted green line uses the result in [8], based on Serfling inequality alone.
@@
On another note, in Appendix IV we include optimal one-sided CIs for the average of independent Bernoulli parameters [30, 31], potentially replacing $\Gamma^{\pm}_{n,\epsilon}(\hat{p})$ (Proposition 1) in the failure probability estimation of decoy-state QKD schemes. These CIs are related to the well-known Clopper-Pearson CIs of the binomial distribution, which in fact have been used in the context of QKD security before (see e.g. [32, 33]).
@@
QKD simulations. Next, we show the impact that the proposed tools have in the finite-key performance of QKD. Let us consider the ideal BBM92 protocol first. Consistently with [15, 8], we evaluate a simplified scenario in which Alice and Bob share a $2N$-partite quantum state as a protocol input. Then, they jointly select between two binary and complementary quantum measurements uniformly at random on a round-by-round basis, in so each of them extracting a raw data block of $N$ bits. These blocks are randomly splitted into a fixed-length test (sifted-key) string of $n$ ($N-n$) bits. That is to say, both the key and the test string are randomly sampled from the two bases. In the simulations, we set the tolerated quantum bit error rate (QBER) to $p^{\rm th}=4.55\%$ for ease of comparison with [15], where this choice is motivated by the results of the Micius experiment [34]. Also, an error correction (EC) leakage of $\lambda_{\rm EC}=1.19(N-n)h(p^{\rm th})$ bits is assumed, $h(x)$ denoting the binary entropy of $x$. The correctness parameter, $\epsilon_{\rm cor}$, and the privacy amplification error, $\epsilon_{\rm PA}$, are both set to $10^{-8}$ for illustration purposes, and the failure probability bound is set to $\epsilon_{\rm PE}=4\times{}10^{-16}$. Altogether, these settings lead to an overall secrecy parameter of $\epsilon_{\rm sec}\approx{}5\times{}10^{-8}$ (see Appendix I).
With these inputs, in Fig. 2a we plot the extractable key length divided by the data block size for blocks up to $N=10^{6}$ bits, optimizing the sampling fraction $n/N$. In addition, we superpose an inset of the minimum block-size regime, showing that our analytical tool (solid blue line) allows to reduce the minimum block size with respect to [15] (dashed orange line) by more than 16%, while our numerical tool (solid red line) attains a reduction beyond 40%. Furthermore, the tool of [15] would roughly require a three-orders-of-magnitude larger $\epsilon_{\rm PE}$ for a nonzero key with the block size collected in the Micius experiment [34], indicated with a dotted red vertical line in the inset. Notably as well, both of our tools perform very similarly —and significantly above [15, 8]— for non-critical block sizes.
@@
Figure 2: Finite secret key rate versus data block size. (a) Ideal BBM92 protocol. The solid red (blue) line corresponds to our numerical (analytical) tool, the dashed orange line uses the result in [15], and the dotted green line uses the result in [8]. The dotted red vertical line in the inset indicates the block size gathered in the Micius experiment [34], $N=3100$ bits. (b) Decoy-state BB84 protocol. The same color code is used again, indicating the random sampling tool under consideration (in particular, the approach of [8] is not presented in this case). The additional confidence intervals required for the decoy-state analysis are selected as specified in the main text. For both figures, the settings employed in the simulations are provided in the main text as well.
@@
Let us now assess the more practical decoy-state BB84 protocol. We consider a scheme with three common intensities per basis, $\mu$ (signal), $\nu$ (decoy) and $\omega$ (vacuum), and employ the $Z$ ($X$) basis for key extraction (PE) with a fixed-size data block of $N_{\rm Z}$ ($N_{\rm X}$) bits drawn at random from the corresponding data pool. Again, in the simulations we contemplate blocks up to $N=N_{\rm Z}+N_{\rm X}=10^{6}$ bits, and superpose an inset of the minimum block-size regime. Naturally, we set $N_{\rm Z}=q_{\rm Z}^{2}/(q_{\rm Z}^{2}+q_{\rm X}^{2})N$, where $q_{\rm Z}$ ($q_{\rm X}$) denotes the $Z$ ($X$) basis probability, common to both parties. Also, while we fix $\omega=10^{-4}$ to account for the finite extinction ratio of the intensity modulator [35], the settings $\mu$ and $\nu$, together with their respective probabilities, $p_{\mu}$ and $p_{\nu}$, and the test basis probability, $q_{\rm X}$, are tuned to optimize the key length per block, $l/N$. For simplicity, in order to select the acceptance thresholds of the PE tests [36], the average behaviour of a typical fiber and detector model is considered (see Appendix II for further details). We assume an overall system loss of 30 dB, a dark count probability of $p_{\rm d}=6\times{}10^{-7}$ [37], and a misalignment error rate of $e_{\rm mis}=5\times 10^{-3}$ [12]. As in the BBM92 protocol, the EC leakage is set to $\lambda_{\rm EC}=1.19N_{\rm Z}h(\theta^{\rm th})$, $\theta^{\rm th}$ denoting the correctable QBER of a key block [38]. Lastly, we set again the overall failure probability to $\epsilon_{\rm PE}=4\times 10^{-16}$, and assume a common value for each individual error term contributing to it [39]. Similarly, we take $\epsilon_{\rm cor}=\epsilon_{\rm PA}=\delta=10^{-8}$, where $\delta>0$ is an intrinsic variable of the security analysis (see Appendix II). Overall, this leads to $\epsilon_{\rm sec}={}7\times{}10^{-8}$.
@@
The results are shown in Fig. 2b. Unlike the BBM92 protocol, the finite-key security of the decoy-state scheme does not rely on the random sampling tool alone, but also on CIs for the average parameter of a set of Bernoulli trials. Indeed, such CIs are repeatedly used in the decoy-state analysis, and thus they have a very critical impact in the performance. As an example, both orange lines in Fig. 2b use the random sampling tool of [15], but the dotted line relies on Hoeffding inequality [17] for the extra CIs (as originally done in [12]), while the dashed one relies on the multiplicative Chernoff bound [20] (following the improvement of [18]). Meanwhile, the solid blue line results from using the random sampling tool of Proposition 2 and the extra CIs of Proposition 1, both originating from our relaxed additive Chernoff bound and each providing an advantage on its own. Lastly, the solid red line uses our numerical upgrades of these two propositions, deploying optimal CIs in both cases. Overall, our analytical tools reduce the minimum block size by more than $22\%$ compared to previous approaches, and the reduction exceeds $38\%$ with the numerical tools. Once more, our analytical and numerical solutions perform very similarly in non-critical regimes, with a strong edge over existing approaches.
@@
Conclusions. The performance of QKD is deeply affected by finite statistics. In satellite QKD [34, 40, 41, 42], the data collection per satellite overpass is limited, and sharp statistics may translate into an enormous saving of time and resources. What is more, one could envision future QKD applications requiring short keys on-demand for daily transactions with high security standards. For these, a poor statistical analysis could result in an unbearable penalty.

### v4

@@ Abstract
The performance of quantum key distribution (QKD) heavily depends on statistical inference. For a broad class of protocols, the central statistical task is a random sampling problem, customarily addressed using a hypergeometric tail bound due to Serfling. Here, we provide an alternative solution for this task of unprecedented tightness among QKD security analyses. As a by-product, confidence intervals for the average of nonidentical Bernoulli parameters follow too. These naturally fit in statistical analyses of decoy-state QKD and also outperform standard tools. Last, we show that, in a vast parameter regime, the use of tail bounds is not enforced because the cumulative mass function of the hypergeometric distribution is accurately computable. This sharply decreases the minimum block sizes necessary for QKD, and reveals the tightness of our analytical bounds when moderate-to-large blocks are considered.
Introduction—Quantum key distribution (QKD) is the only approach to key exchange that is provably secure against computationally unbounded adversaries [1]. Today, considerable efforts are being made to boost the performance of practical QKD systems while guaranteeing their physical-layer security [2, 3]. Among the multiple facets of this ongoing program, it is of utmost importance to achieve tighter lower bounds on the extractable secret key length, a goal that can be tackled by refining QKD security proofs. These proofs build upon quantum information-theoretic results and strongly rely on statistical inference, which plays a pivotal role to keep the necessary data block sizes to a minimum.
To be precise, establishing the finite-key security of a QKD protocol requires to upper bound the probability of a “failure” happening in its parameter estimation (PE) step. This refers to the event where the protocol does not abort but the adversary obtains nontrivial information about the final key [1]. In BB84-like protocols [4], a very popular proof technique is based on an entropic uncertainty relation (EUR) for smooth entropies [5, 6], in which case the simplest formulation of the problem goes as follows. Let us consider a two-piece experiment where a population of $N$ binary random variables is generated according to an arbitrary probability distribution, and a random sample of size $n$ is drawn from the population afterward. If we denote the frequency of ones in the test (complementary) sample as $\hat{p}$ ($\hat{q}$), the goal is to find a threshold value $q^{\rm th}\geq{}p^{\rm th}$ such that
@@
for arbitrary values of $N$, $n$, $\epsilon$ and $p^{\rm th}$. Naturally, an interplay exists between $\epsilon$ and $q^{\rm th}$: fulfilling Eq. (1) with a very small $\epsilon$ comes at the price of tolerating larger values of $q^{\rm th}$. In the context of QKD—in which $\hat{q}$ is often referred to as the phase-error rate—this translates into a direct penalty on the provably secure key length. Notably, the above formulation of the problem shows up exactly in ideal implementations of the BB84 [4] or the BBM92 [7, 8] protocols, and we adhere to the latter in the main text for didactic purposes. The detailed protocol description and secret key length formula can be found in Appendix I. On the other hand, the popular decoy-state BB84 protocol [9, 10, 11, 12], which is implemented in most of today’s commercial QKD setups, is addressed in Appendix II.
@@
In EUR-based security proofs [8, 13], it is commonplace to derive the desired $q^{\rm th}$ by resorting to an exponential tail bound on the hypergeometric (HG) distribution known as Serfling inequality [14]. Recently though, a sharper bound has been derived in [15], combining Serfling inequality with a refined inequality due to Hush and Scovel [16]. Nonetheless, the reported improvement is moderate and heavily relies on the use of very large samples for testing, an undesirable feature given the fact that the final key is extracted from the complementary data.
@@
In this Letter, we propose a simpler approach to calculate $q^{\rm th}$, resorting to a well-known link between random sampling with and without replacement [17]. Remarkably, the bound we obtain outperforms [15] in all the explored parameter regimes relevant for QKD, where one restricts $p^{\rm th}$ to a few percent at most, $N$ to thousands or tens of thousands at least, and $\epsilon$ to rather minuscule values (often below $10^{-10}$). Furthermore, in the derivation of the bound, confidence intervals (CIs) are obtained for the average parameter of a set of independent Bernoulli trials. These intervals, which provide an additional tool required for current decoy-state QKD security proofs, also outperform the analytical results considered for that purpose so far [12, 18]. Last, we show that, contrary to general belief, the cumulative mass function (CMF) of the HG distribution is exactly computable in a vast parameter regime of interest to QKD. Importantly, this enables the calculation of optimal [20, 21, 22] confidence upper bounds for the population parameter following the Clopper-Pearson method [23, 24]. Applying the latter to the random sampling problem of Eq. (1) —and also to its generalization to decoy-state QKD schemes—we report a dramatic decrease of the minimum block sizes necessary for key extraction. On the other hand, explicit comparison with our simple analytical tools reveals the tightness of the latter for moderate-to-large blocks.
@@
Analytical results—Our analytical results are triggered by two observations. First, the additive Chernoff bound [25] (simply Chernoff bound in what follows), conventionally used in the context of sampling with replacement, also holds for sampling without replacement. This feature, which is a consequence of a general theorem by Hoeffding [17]—see also [26] for a simple proof— has already been exploited in QKD security analyses before [27, 28]. Secondly, Serfling inequality—the standard bound used in sampling without replacement—essentially reduces to Hoeffding inequality [17] for small sampling fractions $n/N$. Nevertheless, the latter constitutes a loose relaxation of the Chernoff bound, revealing that this bound must certainly be better for small $n/N$. With this in mind, we derive a tight relaxation of the Chernoff bound and indistinctly use it to obtain one-sided CIs for the population parameters in sampling with and without replacement. Then, we apply these CIs to solve Eq. (1) and its generalization to decoy-state QKD.
@@
A proof of this claim is given in Appendix III. Also, the simpler claims $\Pr\left[\max\{0,\gamma^{-}_{n,\epsilon}(\hat{p})\}>p\right]\leq{}\epsilon$ and $\Pr\left[\min\{1,\gamma^{+}_{n,\epsilon}(\hat{p})\}<p\right]\leq{}\epsilon$ follow identically [30].
Since Proposition 1 is solely derived from the Chernoff bound, it automatically extends to a HG random variable [17, 26]. The upper extension goes as follows.
@@
As shown next, Proposition 2 allows to obtain a threshold value $q^{\rm th}$ fulfilling Eq. (1), which provides the so-called failure probability estimation of the BBM92 protocol.
@@
for the threshold function $q^{\rm th}_{N,n,\epsilon}(x)$ defined in Eq. (6), given that the latter is locally nondecreasing at $p^{\rm th}$.
@@
In the first inequality we use the fact that $\{\hat{p}\leq{}p^{\rm th},\hat{q}\geq{}q^{\rm th}_{N,n,\epsilon}\left(p^{\rm th}\right)\}\implies{}\{\hat{q}\geq{}q^{\rm th}_{N,n,\epsilon}\left(\hat{p}\right)\}$. This is so because $q^{\rm th}_{N,n,\epsilon}(x)$ is nondecreasing for $x\in[0,p^{\rm th}]$ if it is nondecreasing at $x=p^{\rm th}$, and the latter is guaranteed by assumption. The averaging over $\hat{T}$ in the second inequality is a purely ad hoc step to enable the use of Proposition 2, which presumes a fixed ratio of ones in the population. The third inequality follows from Proposition 2. ∎
A relevant observation is that the condition on the slope at $p^{\rm th}$, which is equivalent to $\Gamma^{+^{\prime}}_{n,\epsilon}(p^{\rm th})\geq{}n/N$ [$\Gamma^{+^{\prime}}_{n,\epsilon}(p^{\rm th})$ denoting the derivative of $\Gamma^{+}_{n,\epsilon}(x)$ with respect to $x$ evaluated in $p^{\rm th}$], is unrestrictive for all practical purposes [31].
@@
Numerical results—Certainly, tighter results can be obtained by resorting to numerical methods. For instance, using Stirling’s formula and the Gaussian cumulative density function, the authors of [27] construct a HG tail inequality that outperforms our relaxed Chernoff bound in certain regimes, but does not enable a closed formula for the related CI. Alternatively, one could use the exact Chernoff bound rather than a relaxation. Note, however, that if one is willing to use numerical techniques, tail inequalities may become superfluous, because numerically computing the CMF of the HG distribution is often feasible in QKD applications. Crucially, this enables the calculation of optimal one-sided CIs [22, 20] using the well-known Clopper-Pearson method based on pivoting the CMF [23]. In particular, for our purposes, one can replace $\Gamma^{+}_{n,\epsilon}(\hat{p})$ in Corollary 1 by the tighter statistic $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$, defined via [21, 22, 20]
@@
if the target set of the minimization is nonempty, and via $\mathcal{CP}^{+}_{N,n,\epsilon}(x)=(N+1)/N$ otherwise [32], where we recall that $p$ takes discrete values in the set $\{0,1/N,2/N,\ldots{},1\}$ [33]. Remarkably, accurate determination of the CMF is required in order to compute $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$, with a numerical precision beyond $\epsilon$ according to Eq. (9[34]. Indeed, for a broad range of settings of interest to QKD, this is a manageable task with existing software.
@@
Importantly, since $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$ fulfills Corollary 1, Proposition 2 also holds replacing Eq. (6) by
$q^{\rm th}_{N,n,\epsilon}(x)=\frac{N\mathcal{CP}^{+}_{N,n,\epsilon}(x)-nx}{N-n},$ (10)
and monotonicity arguments allow to further establish Proposition 3 with this alternative threshold function.
@@
In Fig. 1 we compare various candidate threshold values $q^{\rm th}$ satisfying Eq. (1): Eq. (6), Eq. (10), and the thresholds derived in [15] and [8]. For illustration purposes, we consider typical QKD inputs, $p^{\rm th}\leq{}4\%$, $N=10^{5}$, $n=10^{4}$ and $\epsilon=10^{-9}$, and remark that the improvement of our tools—Eq. (6) and Eq. (10)—prevails in all the explored parameter regimes relevant for QKD.
@@
Figure 1: Threshold value $q^{\rm th}$ satisfying Eq. (1), for $p^{\rm th}\leq{}4\%$, $N=10^{5}$, $n=10^{4}$, and $\epsilon=10^{-9}$. The solid red (blue) line corresponds to our numerical (analytical) tool, based on the exact cumulative mass function of the hypergeometric distribution (our relaxed additive Chernoff bound [25]). The dashed orange line uses the result in [15], which combines Serfling [14] and Hush and Scovel [16] inequalities, and the dotted green line uses the result in [8], based on Serfling inequality alone.
@@
Additionally, in Appendix IV we include optimal one-sided CIs for the average of independent Bernoulli parameters [35, 36], potentially replacing $\Gamma^{\pm}_{n,\epsilon}(\hat{p})$ (Proposition 1) in the failure probability estimation of decoy-state QKD schemes. These CIs are related to the Clopper-Pearson CIs for binomial proportions, which in fact have been used in the context of QKD security before (see e.g. [37, 38]).
@@
QKD simulations—Next, we show the impact that the proposed tools have in the finite-key performance of QKD. Let us consider the ideal BBM92 protocol first. Consistently with [15, 8], we evaluate a simplified scenario in which Alice and Bob share a $2N$-partite quantum state as a protocol input. Then, they jointly select between two binary and complementary quantum measurements uniformly at random on a round-by-round basis, each of them extracting a raw data block of $N$ bits. These blocks are randomly split into a fixed-length test (sifted-key) string of $n$ ($N-n$) bits. That is to say, both the key and the test string are randomly sampled from the two bases. In the simulations, we set the tolerated quantum bit error rate (QBER) to $p^{\rm th}=4.55\%$ for ease of comparison with [15], where this choice is motivated by the results of the Micius experiment [39]. Also, an error correction leakage of $\lambda_{\rm EC}=1.19(N-n)h(p^{\rm th})$ bits is assumed, $h(x)$ denoting the binary entropy of $x$. The correctness parameter, $\epsilon_{\rm cor}$, and the privacy amplification error, $\epsilon_{\rm PA}$, are both set to $10^{-8}$ for illustration purposes, and the failure probability bound is set to $\epsilon_{\rm PE}=4\times{}10^{-16}$. Altogether, these settings lead to an overall secrecy parameter of $\epsilon_{\rm sec}=5\times{}10^{-8}$ (see Appendix I).
@@
Figure 2: Finite secret key rate versus data block size. (a) Ideal BBM92 protocol. The dotted green line uses the result in [8], the dashed orange line uses the result in [15] and the solid blue line uses Proposition 3 in this work, originating from our relaxed additive Chernoff bound. Last, the solid red line corresponds to the numerical upgrade of Proposition 3, which replaces the threshold function of Eq. (6)—relaxed-Chernoff—by that of Eq. (10)—Clopper-Pearson—. The dotted black vertical line in the inset indicates the block size gathered in the Micius experiment [39], $N=3100$ bits. (b) Decoy-state BB84 protocol. Both orange lines use the random sampling tool of [15], but the dotted (dashed) line relies on Hoeffding’s inequality [17] (the tighter multiplicative Chernoff bound deployed in [18]) for the extra confidence intervals for Bernoulli trials required by the decoy-state method. The solid blue line uses the random sampling tool of Proposition 2 and the extra confidence intervals of Proposition 1, both based on our relaxed Chernoff bound. Finally, the solid red line uses our numerical upgrades of these two propositions, deploying optimal confidence intervals in both cases [20, 35]. The protocol inputs and experimental settings considered in the simulations are listed in the main text.
With these inputs, in Fig. 2(a) we plot the extractable key length divided by the data block size for blocks up to $N=10^{6}$ bits, optimizing the sampling fraction $n/N$. In addition, we superpose an inset of the minimum block-size regime, showing that our analytical tool (solid blue line) reduces the minimum block size with respect to [15] (dashed orange line) by more than 16%, while our numerical tool (solid red line) attains a reduction beyond 40%. Furthermore, the tool of [15] would roughly require a 3-orders-of-magnitude larger $\epsilon_{\rm PE}$ for a nonzero key with the block size collected in the Micius experiment [39] (dotted black vertical line in the inset). Notably as well, both of our tools perform very similarly—and significantly above [15, 8]—for noncritical block sizes.
@@
Let us now assess the more practical decoy-state BB84 protocol. We consider a scheme with three common intensities per basis, $\mu$ (signal), $\nu$ (decoy) and $\omega$ (vacuum), and employ the $Z$ ($X$) basis for key extraction (PE) with a fixed-size data block of $N_{\rm Z}$ ($N_{\rm X}$) bits drawn at random from the corresponding data pool. Again, in the simulations we contemplate blocks up to $N=N_{\rm Z}+N_{\rm X}=10^{6}$ bits, and superpose an inset of the minimum block-size regime. Naturally, we set $N_{\rm Z}=q_{\rm Z}^{2}/(q_{\rm Z}^{2}+q_{\rm X}^{2})N$, where $q_{\rm Z}$ ($q_{\rm X}$) denotes the $Z$ ($X$) basis probability, common to both parties. Also, we fix $\omega=10^{-4}$ to account for the finite extinction ratio of the intensity modulator [40]. The settings $\mu$ and $\nu$, together with their respective probabilities, $p_{\mu}$ and $p_{\nu}$, and the test basis probability, $q_{\rm X}$, are tuned to optimize the key length per block, $l/N$. For simplicity, in order to select the acceptance thresholds of the PE tests [41], the average behavior of a typical fiber and detector model is considered (see Appendix II for the details). We assume an overall system loss of 30 dB, a dark count probability of $p_{\rm d}=6\times{}10^{-7}$ [42], and a misalignment error rate of $e_{\rm mis}=5\times 10^{-3}$ [12]. As in the BBM92 protocol, the error correction leakage is set to $\lambda_{\rm EC}=1.19N_{\rm Z}h(\theta^{\rm th})$, $\theta^{\rm th}$ denoting the correctable QBER of a key block [43]. Last, we set again the overall failure probability to $\epsilon_{\rm PE}=4\times 10^{-16}$, and assume a common value for each individual error term contributing to it [44]. Similarly, we take $\epsilon_{\rm cor}=\epsilon_{\rm PA}=\delta=10^{-8}$, where $\delta>0$ is an intrinsic variable of the security analysis (see Appendix II). Overall, this leads to $\epsilon_{\rm sec}={}7\times{}10^{-8}$.
@@
The results are shown in Fig. 2(b). Unlike the BBM92 protocol, the security of the decoy-state scheme does not rely on the random sampling tool alone, but also on CIs for the average parameter of a set of Bernoulli trials. In fact, such CIs are repeatedly used in the decoy-state analysis, and thus they have a critical impact in the performance. The reader is referred to the caption for the specifications of the different lines in Fig. 2(b). Overall, our analytical tools reduce the minimum block size by more than $22\%$ compared to previous approaches, and the reduction exceeds $38\%$ with the numerical tools. Once more, our analytical and numerical solutions perform very similarly in noncritical regimes, with a strong edge over existing approaches.
On top of all the above results, in Appendices V and VI we survey analytical HG bounds available in the literature, and derive two tools that perform better than our Chernoff-based confidence upper bound (Corollary 1) for small blocks. Particularly, in this regime, we obtain a performance comparable to the numerical one by reproducing our analytical method but with Hush and Scovel inequality.
@@
Conclusions—The performance of QKD is deeply affected by finite statistics. In satellite QKD [39, 45, 46, 47], the data collection per satellite overpass is limited, and sharp statistics may entail an enormous saving of time and resources. Moreover, one could envision future applications requiring short keys ondemand for daily transactions with high security standards. For these, a poor statistical analysis would result in an unbearable penalty.